the rate constant for a first order reaction is 2.3*10^-4 . Calculate the time taken to complete 2/3 of the reaction?
 

Grade:12th pass

1 Answers

User
5 years ago
For first order reaction
t=2.303/2.3*10^-4[log (a/a-x)]
 =2.303/2.3*10^-4[log(1/1-2/3)]
  =2.303/2.3*10^-4(log3)
 =4777.436 units